find P(-2.31 < Z ≤ 0.82).


1
Expert's answer
2021-09-27T23:55:02-0400
"P(-2.31<Z\leq0.82)=P(Z\leq 0.82)-P(Z\leq-2.31)"

"\approx0.7938919-0.0104441\approx0.783448"

"P(-2.31<Z\leq0.82)=0.783448"